OpenStudy (anonymous):

Four pine seedling and 3 fir seedlings cost 29.00. Two pine seedlings and 7 fir seedlings cost $42.00 Determine the cost of each type of seedling.

hero (hero):

4p + 3f = 29 2p + 7f = 42

hero (hero):

Solve the system

OpenStudy (anonymous):

I tried but I got a weird answer? Do you mind me writing what I did and tell me where I went wrong?

OpenStudy (anonymous):

@Hero

hero (hero):

I don't know where you went wrong because you haven't posted your work

OpenStudy (anonymous):

I am, hold on.

OpenStudy (anonymous):

4x+3y=29 2x+7y=42 ___________ 6x+10y=29 ... 6x+10y=29 -6x -6x 10y=29-6x /10 /10 y=2.9-0.6x ... 4x+3(2.9-0.6x)=29 4x+8.7+2.04=29 Have I done anything wrong so far?

OpenStudy (anonymous):

@Hero

hero (hero):

Who instructed you to add them together. You are not authorized to simply add them together.

OpenStudy (anonymous):

Can you show me? I don't get the difference?

OpenStudy (anonymous):

@Hero